OpenStudy (gio1994):

Calculus 3!!!! Help 1. Compute the line integral of the vector field F= < y2+2-cox(x4) , 3+sin (y6) +4xy> , for the closed curve consisting of the x-axis from ( 0 , 0 ) to ( 3 , 0 ) , followed by the quarter circle (centered at the origin) from ( 3 , 0 ) to ( 0 , 3 ), and back to ( 0 , 0 ) along the y-axis. 2. For the solid S bounded by x= y2+ z2 and x = 9 , find: a) the volume; b) the centroid.

OpenStudy (518nad):

wheres the vec field

OpenStudy (gio1994):

Compute the line integral of the vector field F= < y2+2-cox(x4) , 3+sin (y6) +4xy> , for the closed curve consisting of the x-axis from ( 0 , 0 ) to ( 3 , 0 ) , followed by the quarter circle (centered at the origin) from ( 3 , 0 ) to ( 0 , 3 ), and back to ( 0 , 0 ) along the y-axis. For the solid S bounded by x= y2+ z2 and x = 9 , find: a) the volume; b) the centroid.

OpenStudy (gio1994):

The Question did not copy over the first time sorry

OpenStudy (518nad):

parametrize the curve ur integrating over

OpenStudy (518nad):

we will get it in this form F(x(t),y(t)) dot r'(t)/|r'(t)| dt

OpenStudy (518nad):

r(t) is our parametrized curve and we will rewrite our vector field F in terms of our parameter

OpenStudy (518nad):

F.dR dR=dR/dt * dt

OpenStudy (518nad):

go ahead and parametrize, or find the bounds a circle with radius 1 x=cost , y= sin t a horizontal line on xaxis is of form x=t,y=0 use this to figure out your curve

OpenStudy (gio1994):

Okay. I will try that

ganeshie8 (ganeshie8):

Green's theorem, if you have covered it already, works nicely here..

OpenStudy (holsteremission):

Green's seems to be the only choice. Integrating along the horizontal part would be difficult without resorting to numerical methods: \[\int_{C_1}\mathbf F\cdot\mathrm d\vec{r}_1=\int_0^1\left\langle2-\cos(81 t^4),3\right\rangle\cdot\left\langle3,0\right\rangle\,\mathrm dt=3\int_0^1(2-\cos(81t^4))\,\mathrm dt\]There is an antiderivative in terms of the incomplete gamma function, but I doubt that would be covered in a Calc III course. http://mathworld.wolfram.com/IncompleteGammaFunction.html If you write \(\mathbf F(x,y)=\langle P(x,y),Q(x,y)\rangle\), then by Green's theorem you have \[\int_C\mathbf F\cdot\mathrm d\vec{r}=\iint_D\left(\frac{\partial Q}{\partial x}-\frac{\partial P}{\partial y}\right)\,\mathrm dx\,\mathrm dy\]where \(D\) is the region bounded by \(C\). With \(\mathbf F=\left\langle \underbrace{y^2+2-\cos x^4}_P,\underbrace{3+\sin y^6+4xy}_Q\right\rangle\), you have \[\int_C\mathbf F\cdot\mathrm d\vec{r}=\iint_D(4y-2y)\,\mathrm dx\,\mathrm dy=2\iint_Dy\,\mathrm dx\,\mathrm dy\]which can be computed easily if you convert to polar coordinates.